OpenStudy (anonymous):

value of K to make a perfect square: p+1/4root(p)+k^2?

hartnn (hartnn):

hey aryan :) \(for ~ ~ x^2+bx\) the constant that can be added to make the expression, a perfect square is \((b/2)^2 \) so for \(p +(1/4)\sqrt p \) we add, \(\Large k^2 = (1/2 \times 1/4)^2\) find k from here :)

hartnn (hartnn):

basically just half of the co-efficient of middle term :)

OpenStudy (anonymous):

yeah got it. Answer is +/-(1/8).

hartnn (hartnn):

thats \(\huge \checkmark \) :)

OpenStudy (anonymous):

Thanks for helping :)

hartnn (hartnn):

welcome ^_^